For the equilibrium: 2 SO3(g) < = > O2(g) + 2 SO2(g) Kp = 0.269 at 625 oC


What is Kc at this temperature? Kp = Kc[RT]Δn R = 0.08206 L-atm/mol K
Answer #1

26039) 029)+ 2 802,69) Kp=0.269 on = no of ganeous products - no. of gaseous reactants E 3-2 R=0008206 L.atm/molik T= 625+273